calculate ΔG at a) 20oC, b) 100oC and c) 150oC.

1
Expert's answer
2022-05-02T12:51:02-0400

ΔG=ΔH-293ΔS, ΔG=ΔH-373ΔS, ΔG=ΔH-423ΔS